OpenStudy (anonymous):

What is the equation of the line that passes through the point (2, 5) and has a slope of 2?

OpenStudy (anonymous):

no

Mehek (mehek14):

\(y=\color{blue}{m}x+\color{red}{b}\)

Mehek (mehek14):

where \(\color{blue}{m}=slope\\ \color{red}{b}=y-intercept\)

Mehek (mehek14):

you have slope = 2 so plug that into the equation \(y=2x+\color{red}{b}\)

OpenStudy (anonymous):

alright

Mehek (mehek14):

we need to find b plug in the ordered pair (2,5)

Mehek (mehek14):

\(\large{x = 2\\y=5}\)

OpenStudy (anonymous):

y= 4 + 5?

Mehek (mehek14):

\(\bf{\Large{5=2(2)+b}}\)

Mehek (mehek14):

you put 5 instead of y

OpenStudy (anonymous):

oh alright

Mehek (mehek14):

2 * 2 = 4 \(\bf{5=4+b}\)

OpenStudy (anonymous):

b=1?

Mehek (mehek14):

yes b = 1

Mehek (mehek14):

so now what's the equation?

OpenStudy (anonymous):

5= 4 +1

Mehek (mehek14):

no in \(\bf{y=mx+b}\)

Mehek (mehek14):

m = 2 b = 1

OpenStudy (anonymous):

y = 4 + 1....?

Mehek (mehek14):

no... in \(\bf{y=\color{blue}{m}x+\color{red}{b}}\) put m = 2 and b = 1

OpenStudy (anonymous):

y= 2x + 1

Mehek (mehek14):

and that's your answer ^_^

OpenStudy (anonymous):

god i just realize how far i fell behind haha

OpenStudy (anonymous):

thank you

Mehek (mehek14):

yw ^_^